### v1

@@ Appendix F Proofs of Extensions (Section 6)
      The POCS update is equivalent to an SGD step: $\mathbf{w}_{t}={\Pi}_{\tau(t)}(\mathbf{w}_{t-1})=\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})$.
@@
First, by Theorem 1.5.5 in Facchinei and Pang (2003), $f_{m}$ is continuously differentiable and for every $\mathbf{w}\in\mathbb{R}^{d},m\in\left[{T}\right]$, $\nabla f_{m}(\mathbf{w})=\mathbf{w}-{\Pi}_{m}(\mathbf{w})$. Plugging in $\nabla f_{\tau(t)}(\mathbf{w}_{t-1})$ into an appropriate SGD step, we get
$\mathbf{w}_{t}=\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})=\mathbf{w}_{t-1}-\left(\mathbf{w}_{t-1}-{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\right)={\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\,,$
and the second part of the lemma follows. In addition, $\forall\mathbf{x},\mathbf{w}\in\mathbb{R}^{d}$, we prove convexity by using a projection inequality (also from Theorem 1.5.5 in Facchinei and Pang, 2003). That is,
$f_{m}(\mathbf{x})-f_{m}(\mathbf{w})-\langle\nabla f_{m}(\mathbf{w}),\mathbf{x}-\mathbf{w}\rangle=\frac{1}{2}\|\mathbf{x}-{\Pi}_{m}(\mathbf{x})\|^{2}-\frac{1}{2}\|\mathbf{w}-{\Pi}_{m}(\mathbf{w})\|^{2}-\langle\mathbf{w}-{\Pi}_{m}(\mathbf{w}),\mathbf{x}-\mathbf{w}\rangle$
$\hskip 41.00012pt=\frac{1}{2}\|\mathbf{x}-{\Pi}_{m}(\mathbf{x})\|^{2}-\frac{1}{2}\|\mathbf{w}-{\Pi}_{m}(\mathbf{w})\|^{2}-\langle\mathbf{w}-{\Pi}_{m}(\mathbf{w}),\mathbf{x}-{\Pi}_{m}(\mathbf{x})\rangle$
$\hskip 75.0001pt+\langle\mathbf{w}-{\Pi}_{m}(\mathbf{w}),{\Pi}_{m}(\mathbf{w})-{\Pi}_{m}(\mathbf{x})\rangle+\langle\mathbf{w}-{\Pi}_{m}(\mathbf{w}),\mathbf{w}-{\Pi}_{m}(\mathbf{w})\rangle$
$\hskip 41.00012pt\geq\frac{1}{2}\|\mathbf{x}-{\Pi}_{m}(\mathbf{x})\|^{2}-\frac{1}{2}\|\mathbf{w}-{\Pi}_{m}(\mathbf{w})\|^{2}-\langle\mathbf{w}-{\Pi}_{m}(\mathbf{w}),\mathbf{x}-{\Pi}_{m}(\mathbf{x})\rangle+0+\|\mathbf{w}-{\Pi}_{m}(\mathbf{w})\|^{2}$
$\hskip 41.00012pt=\frac{1}{2}\|\mathbf{x}-{\Pi}_{m}(\mathbf{x})-\mathbf{w}+{\Pi}_{m}(\mathbf{w})\|^{2}\geq 0\,.$
@@
$\left\|{\nabla f_{m}(\mathbf{x})-\nabla f_{m}(\mathbf{w})}\right\|=\left\|{\mathbf{x}-{\Pi}_{m}(\mathbf{x})-\left({\mathbf{w}-{\Pi}_{m}(\mathbf{w})}\right)}\right\|=\left\|{(\mathbf{I}-{\Pi}_{m})(\mathbf{x})-(\mathbf{I}-{\Pi}_{m})(\mathbf{w})}\right\|\leq\left\|{\mathbf{x}-\mathbf{w}}\right\|\,,$
where we used the non-expansiveness of $\mathbf{I}-{\Pi}_{m}$ (Propositions 4.2, 4.8 in Bauschke et al., 2011).
Lemma 37.
Let $\mathcal{K}\subseteq\mathbb{R}^{d}$ be a nonempty closed and convex set, and $f(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w})}\right\|^{2}$. Then, we have for any $\mathbf{z}\in\mathbb{R}^{d}$ and $\gamma>0$
@@
We have $\nabla f(\mathbf{w})=\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w})$. Hence, by Theorem 1.5.5 in Facchinei and Pang (2003),
$\left\langle\nabla f(\mathbf{w}),\mathbf{w}-\mathbf{z}\right\rangle$ $=\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{w}-\mathbf{z}\right\rangle=\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w})\right\rangle+\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),{\Pi}_{\mathcal{K}}(\mathbf{w})-\mathbf{z}\right\rangle$
$=2f(\mathbf{w})+\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),{\Pi}_{\mathcal{K}}(\mathbf{w})-{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle-\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle$
$\geq 2f(\mathbf{w})-\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle\,.$
Plugging in $\mathbf{z}=\mathbf{u}$, the second term vanishes (since $\mathbf{u}-{\Pi}_{\mathcal{K}}(\mathbf{u})=\mathbf{0}$) and the second claim follows.
@@
$\left\langle\nabla f(\mathbf{w}),\mathbf{w}-\mathbf{z}\right\rangle$ $=2f(\mathbf{w})-\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle$
$\geq 2f(\mathbf{w})-\frac{\gamma}{2}\left\|{\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w})}\right\|^{2}-\frac{1}{2\gamma}\left\|{\mathbf{z}-{\Pi}_{\mathcal{K}}(\mathbf{z})}\right\|^{2}=2f(\mathbf{w})-\gamma f(\mathbf{w})-\frac{1}{\gamma}f(\mathbf{z})\,.$
  • Recall Theorem 14.Consider the same conditions of Reduction 3 and assume a nonempty set intersection $\mathcal{C}_{\star}=\bigcap_{m=1}^{T}\mathcal{C}_{m}\neq\varnothing$. Then, under a random ordering with or without replacement, the expected “residual” of Scheme 4 after $\forall k\geq 1$ iterations (without replacement: $k\in\left[{T}\right]$) is bounded as,
    $\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}{\sum}_{m=1}^{T}\left\|{\mathbf{w}_{k}-{\Pi}_{m}(\mathbf{w}_{k})}\right\|^{2}\Big{]}=\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}{\sum}_{m=1}^{T}\mathrm{dist}^{2}(\mathbf{w}_{k},\mathcal{C}_{m})\Big{]}\leq\frac{7}{\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$
@@
    The proof largely follows the same steps of Theorems 8 and 9. Let $\tau$ be any random ordering, $\mathbf{w}_{0}\in\mathbb{R}^{d}$ an initialization, and $\mathbf{w}_{1},\ldots,\mathbf{w}_{k}$ be the corresponding iterates produced by Scheme 4. By Reduction 3, these are exactly the (stochastic) gradient descent iterates produced when initializing at $\mathbf{w}_{0}$ and using a step size of $\eta=1$, on the $1$-smooth loss sequence $f_{\tau(1)},\ldots,f_{\tau(k)}$ defined by:
    $f_{m}(\mathbf{w})\triangleq\frac{1}{2}\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w}))}\right\|^{2}.$
@@
    $\bar{f}(\mathbf{w})\triangleq\mathbb{E}_{m\sim\textrm{Unif}([T])}f_{m}(\mathbf{w})=\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w})}\right\|^{2}\,.$
    Now, for a with-replacement ordering $\tau$, invoke Theorem 10, except we use Lemma 37 in the proof instead of Lemma 32, to obtain:
@@
For a without-replacement ordering $\tau$, invoke Theorem 13 (with $\eta=1/\beta$), except again we use Lemma 37 in the proof instead of Lemma 32, to obtain:
@@
Similarly, by Lemma 33,
$\mathbb{E}_{\tau}f_{\tau(k+1)}(\mathbf{w}_{k})\triangleq\mathbb{E}_{\tau}\tfrac{1}{2}\left\|\mathbf{w}_{k}-{\Pi}_{\tau\left(k+1\right)}\left(\mathbf{w}_{k}\right)\right\|^{2}\leq\frac{e}{2\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$ ($\tau$ without-replacement)
Combining the last two displays with Proposition 17, we now obtain:
$\mathbb{E}_{\tau}\bar{f}(\mathbf{w}_{k})$ $\triangleq\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{w}_{k}-{\Pi}_{m}(\mathbf{w}_{k})}\right\|^{2}\Big{]}$ ($\tau$ without-replacement)
$=\frac{k}{T}\mathbb{E}_{\tau}\hat{f}_{0:k-1}(\mathbf{w}_{k})+\frac{T-k}{2T}\mathbb{E}_{\tau}\left\|\mathbf{w}_{k}-{\Pi}_{\tau\left(k+1\right)}\left(\mathbf{w}_{k}\right)\right\|^{2}$
@@
which proves the without-replacement case and thus completes the proof.
  • Recall Theorem 15.Under a random ordering, with or without replacement, over $T$ jointly separable tasks, the expected forgetting of the weakly-regularized Scheme 5 (at $\lambda\to 0$) after $k\geq 1$ iterations (without replacement: $k\in\left[{T}\right]$) is bounded as
@@
    $f_{m}(\mathbf{w})$ $\triangleq\frac{1}{2}\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w}))}\right\|^{2}$
    $\bar{f}(\mathbf{w})$ $\triangleq\mathbb{E}_{m\sim\textrm{Unif}([T])}f_{m}(\mathbf{w})=\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w})}\right\|^{2}\,.$
    For $\tau$ sampled with replacement, by Lemma 38 (given below) and the with-replacement result (inside the proof) of Theorem 14, we have
@@
    For $\tau$ sampled without replacement, as argued in Theorem 14, by Lemma 33:
@@
    which completes the proof.
Lemma 38.
@@
Our proof here mostly follows the proof of Lemma 34. Recall that from Eq. (8), any $\beta$-smooth realizable function $h:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ holds that $\left|h(\tilde{\mathbf{w}})-h(\mathbf{w})\right|\leq h(\mathbf{w})+\beta\left\|{\tilde{\mathbf{w}}-\mathbf{w}}\right\|^{2}$. Denote $f_{t}\triangleq f(\cdot;i_{t})$ for all $t\in\left\{0,.\hskip 0.8pt\!.\hskip 0.8pt\!.,T\right\}$. Now, by the standard stability $\iff$ generalization argument (Shalev-Shwartz et al., 2010; Hardt et al., 2016), and denoting by $\mathbf{w}_{\tau}^{(i)}$ the SGD iterate after $\tau$ steps on the training set where the $i$th example was resampled as $j_{i}$:
@@
Next, we bound $\left\|{\mathbf{w}_{T}^{(i)}-\mathbf{w}_{T}}\right\|^{2}$. By the non-expansiveness of gradient steps in the convex and $\beta$-smooth regime when $\eta\leq 2/\beta$ (see Lemma 3.6 in Hardt et al., 2016):
@@
Finally, by the regret bound given in Lemma 11, i.e., $\sum_{i=0}^{T-1}f_{i}(\mathbf{w}_{i})\leq\frac{\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta(2-\eta\beta)}$, we have

### v2

@@ Appendix F Supplementary Material for the Extension Section (Section 6)
      The POCS update is equivalent to an SGD step: $\mathbf{w}_{t}=\bm{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})=\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})$.
@@
    First, by Theorem 1.5.5 in Facchinei and Pang [17], $f_{m}$ is continuously differentiable and for every $\mathbf{w}\in\mathbb{R}^{d},m\in\left[{T}\right]$, $\nabla f_{m}(\mathbf{w})=\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})$. Plugging in $\nabla f_{\tau(t)}(\mathbf{w}_{t-1})$ into an appropriate SGD step, we get
    $\mathbf{w}_{t}=\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})=\mathbf{w}_{t-1}-\left(\mathbf{w}_{t-1}-\bm{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\right)=\bm{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\,,$
    and the second part of the lemma follows. In addition, $\forall\mathbf{x},\mathbf{w}\in\mathbb{R}^{d}$, we prove convexity by using a projection inequality (also from Theorem 1.5.5 in 17). That is,
    $f_{m}(\mathbf{x})-f_{m}(\mathbf{w})-\langle\nabla f_{m}(\mathbf{w}),\mathbf{x}-\mathbf{w}\rangle$
    $=\frac{1}{2}\|\mathbf{x}-\bm{\Pi}_{m}(\mathbf{x})\|^{2}-\frac{1}{2}\|\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})\|^{2}-\langle\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w}),\mathbf{x}-\mathbf{w}\rangle$
    $=\frac{1}{2}\|\mathbf{x}-\bm{\Pi}_{m}(\mathbf{x})\|^{2}-\frac{1}{2}\|\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})\|^{2}-\langle\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w}),\mathbf{x}-\bm{\Pi}_{m}(\mathbf{x})\rangle$
    $\hskip 35.00005pt+\langle\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w}),\bm{\Pi}_{m}(\mathbf{w})-\bm{\Pi}_{m}(\mathbf{x})\rangle+\langle\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w}),\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})\rangle$
    $\geq\frac{1}{2}\|\mathbf{x}-\bm{\Pi}_{m}(\mathbf{x})\|^{2}-\frac{1}{2}\|\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})\|^{2}-\langle\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w}),\mathbf{x}-\bm{\Pi}_{m}(\mathbf{x})\rangle+0+\|\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})\|^{2}$
    $=\frac{1}{2}\|\mathbf{x}-\bm{\Pi}_{m}(\mathbf{x})-\mathbf{w}+\bm{\Pi}_{m}(\mathbf{w})\|^{2}\geq 0\,.$
@@
    $\left\|{\nabla f_{m}(\mathbf{x})-\nabla f_{m}(\mathbf{w})}\right\|$ $=\left\|{\mathbf{x}-\bm{\Pi}_{m}(\mathbf{x})-\left({\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})}\right)}\right\|$
    $=\left\|{(\mathbf{I}-\bm{\Pi}_{m})(\mathbf{x})-(\mathbf{I}-\bm{\Pi}_{m})(\mathbf{w})}\right\|\leq\left\|{\mathbf{x}-\mathbf{w}}\right\|\,,$
    where we used the non-expansiveness of $\mathbf{I}-\bm{\Pi}_{m}$ [Propositions 4.2, 4.8 in 3]. $\blacksquare$
Lemma F.1.
Let $\mathcal{K}\subseteq\mathbb{R}^{d}$ be a nonempty closed and convex set, and $f(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w})}\right\|^{2}$. Then, we have for any $\mathbf{z}\in\mathbb{R}^{d}$ and $\gamma>0$
@@
    We have $\nabla f(\mathbf{w})=\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w})$. Hence, by Theorem 1.5.5 in Facchinei and Pang [17],
    $\left\langle\nabla f(\mathbf{w}),\mathbf{w}-\mathbf{z}\right\rangle$ $=\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{w}-\mathbf{z}\right\rangle$
    $=\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w})\right\rangle+\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\bm{\Pi}_{\mathcal{K}}(\mathbf{w})-\mathbf{z}\right\rangle$
    $=2f(\mathbf{w})+\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\bm{\Pi}_{\mathcal{K}}(\mathbf{w})-\bm{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle-\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-\bm{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle$
    $\geq 2f(\mathbf{w})-\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-\bm{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle\,.$
    Plugging in $\mathbf{z}=\mathbf{u}$, the second term vanishes (since $\mathbf{u}-\bm{\Pi}_{\mathcal{K}}(\mathbf{u})=\mathbf{0}$) and the second claim follows.
@@
    $\left\langle\nabla f(\mathbf{w}),\mathbf{w}-\mathbf{z}\right\rangle$ $=2f(\mathbf{w})-\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-\bm{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle$
    $\geq 2f(\mathbf{w})-\frac{\gamma}{2}\left\|{\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w})}\right\|^{2}-\frac{1}{2\gamma}\left\|{\mathbf{z}-\bm{\Pi}_{\mathcal{K}}(\mathbf{z})}\right\|^{2}$
    $=2f(\mathbf{w})-\gamma f(\mathbf{w})-\frac{1}{\gamma}f(\mathbf{z})\,.$
    $\blacksquare$
  • Recall Theorem 6.1.Consider the same conditions of Reduction 3 and assume a nonempty set intersection $\mathcal{C}_{\star}=\bigcap_{m=1}^{T}\mathcal{C}_{m}\neq\varnothing$. Then, under a random ordering with or without replacement, the expected “residual” of Scheme 4 after $\forall k\geq 1$ iterations (without replacement: $k\in\left[{T}\right]$) is bounded as,
    $\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}{\sum}_{m=1}^{T}\left\|{\mathbf{w}_{k}-\bm{\Pi}_{m}(\mathbf{w}_{k})}\right\|^{2}\Big{]}=\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}{\sum}_{m=1}^{T}\mathrm{dist}^{2}(\mathbf{w}_{k},\mathcal{C}_{m})\Big{]}\leq\frac{7}{\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$
@@
      The proof largely follows the same steps of Theorems 4.2 and 4.3. Let $\tau$ be any random ordering, $\mathbf{w}_{0}\in\mathbb{R}^{d}$ an initialization, and $\mathbf{w}_{1},\ldots,\mathbf{w}_{k}$ be the corresponding iterates produced by Scheme 4. By Reduction 3, these are exactly the (stochastic) gradient descent iterates produced when initializing at $\mathbf{w}_{0}$ and using a step size of $\eta=1$, on the $1$-smooth loss sequence $f_{\tau(1)},\ldots,f_{\tau(k)}$ defined by:
      $f_{m}(\mathbf{w})\triangleq\frac{1}{2}\left\|{\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w}))}\right\|^{2}.$
@@
      $\bar{f}(\mathbf{w})\triangleq\mathbb{E}_{m\sim\textrm{Unif}([T])}f_{m}(\mathbf{w})=\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})}\right\|^{2}\,.$
      Now, for a with-replacement ordering $\tau$, invoke Theorem 5.1, except we use Lemma F.1 in the proof instead of Lemma E.3, to obtain:
@@
    For a without-replacement ordering $\tau$, invoke Theorem E.4 (with $\eta=1/\beta$), except again we use Lemma F.1 in the proof instead of Lemma E.3, to obtain:
@@
    Similarly, by Lemma E.5,
    $\mathbb{E}_{\tau}f_{\tau(k+1)}(\mathbf{w}_{k})\triangleq\mathbb{E}_{\tau}\tfrac{1}{2}\left\|\mathbf{w}_{k}-\bm{\Pi}_{\tau\left(k+1\right)}\left(\mathbf{w}_{k}\right)\right\|^{2}\leq\frac{e}{2\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$ ($\tau$ without-replacement)
    Combining the last two displays with Proposition B.2, we now obtain:
    $\mathbb{E}_{\tau}\bar{f}(\mathbf{w}_{k})$ $\triangleq\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{w}_{k}-\bm{\Pi}_{m}(\mathbf{w}_{k})}\right\|^{2}\Big{]}$ ($\tau$ without-replacement)
    $=\frac{k}{T}\mathbb{E}_{\tau}\hat{f}_{0:k-1}(\mathbf{w}_{k})+\frac{T-k}{2T}\mathbb{E}_{\tau}\left\|\mathbf{w}_{k}-\bm{\Pi}_{\tau\left(k+1\right)}\left(\mathbf{w}_{k}\right)\right\|^{2}$
@@
    which proves the without-replacement case and thus completes the proof. $\blacksquare$
  • Recall Theorem 6.2.Under a random ordering, with or without replacement, over $T$ jointly separable tasks, the expected forgetting of the weakly-regularized Scheme 5 (at $\lambda\to 0$) after $k\geq 1$ iterations (without replacement: $k\in\left[{T}\right]$) is bounded as
@@
    $f_{m}(\mathbf{w})$ $\triangleq\frac{1}{2}\left\|{\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w}))}\right\|^{2}$
    $\bar{f}(\mathbf{w})$ $\triangleq\mathbb{E}_{m\sim\textrm{Unif}([T])}f_{m}(\mathbf{w})=\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})}\right\|^{2}\,.$
    For $\tau$ sampled with replacement, by Lemma F.2 (given below) and the with-replacement result (inside the proof) of Theorem 6.1, we have
    $\mathbb{E}_{\tau}[F_{\tau}(k)]=\mathbb{E}\hat{f}_{0:k-1}(\mathbf{w}_{k})$ $\leq 2\mathbb{E}\bar{f}(\mathbf{w}_{k})+\frac{4\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{k}$
@@
    For $\tau$ sampled without replacement, as argued in Theorem 6.1, by Lemma E.5:
@@
    which completes the proof. $\blacksquare$
Lemma F.2.
@@
    Our proof here mostly follows the proof of Lemma E.6. Recall that from Eq. (8), any $\beta$-smooth realizable function $h:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ holds that $\left|h(\tilde{\mathbf{w}})-h(\mathbf{w})\right|\leq h(\mathbf{w})+\beta\left\|{\tilde{\mathbf{w}}-\mathbf{w}}\right\|^{2}$. Denote $f_{t}\triangleq f(\cdot;i_{t})$ for all $t\in\left\{0,.\hskip 0.8pt\!.\hskip 0.8pt\!.,T\right\}$. Now, by the standard stability $\iff$ generalization argument [61, 25], and denoting by $\mathbf{w}_{\tau}^{(i)}$ the SGD iterate after $\tau$ steps on the training set where the $i$th example was resampled as $j_{i}$:
@@
    $=\mathbb{E}\bar{f}(\mathbf{w}_{T})+\frac{\beta}{T}\sum_{i=0}^{T-1}\mathbb{E}\left\|{\mathbf{w}_{T}^{(i)}-\mathbf{w}_{T}}\right\|^{2}.$
    Next, we bound $\left\|{\mathbf{w}_{T}^{(i)}-\mathbf{w}_{T}}\right\|^{2}$. By the non-expansiveness of gradient steps in the convex and $\beta$-smooth regime when $\eta\leq 2/\beta$ [see Lemma 3.6 in 25]:
@@
    Finally, by the regret bound given in Lemma 5.2, i.e., $\sum_{i=0}^{T-1}f_{i}(\mathbf{w}_{i})\leq\frac{\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta(2-\eta\beta)}$, we have